A slingshot is a child's toy that converts the strain energy of the stretched material into the moving energy of the propelled object. A boy pulls on his slingshot with a force of20 N and it extends by 7 cm. At what velocity does the 20 g stone fly off?


1
Expert's answer
2021-07-28T00:39:01-0400

"A=E,"

"Fs=\frac{mv^2}2,"

"v=\sqrt{\frac{2Fs}m},"

"v=11.8~\frac ms."